OpenStudy (anonymous):

Annie wrote the equation shown below. 3(3y) = (3 • 3) y Which property did Annie apply? associative property of multiplication commutative property of multiplication multiplication property of equality distributive property of multiplication

OpenStudy (anonymous):

multiplication property of equality

OpenStudy (unklerhaukus):

associative property of multiplication

OpenStudy (anonymous):

^ what are those terms? can someone explain to me?

OpenStudy (lgbasallote):

should i say another option? commutative property

OpenStudy (anonymous):

I think @UnkleRhaukus was correct.

OpenStudy (unklerhaukus):

Associativity \[A\star (B\star C)=(A\star B)\star C\] Commutativity \[A\star B=B\star A\] Distributitivity \[A\cdot(B\star C)=(A\cdot B)\star(A\cdot B)\]